Determine the magnitudes of the following vectors;

(i) 𝑨⃗⃗ =(πŸπ’ŠΜ‚+πŸ‘π’‹Μ‚+πŸ”π’ŒΜ‚)

(ii) 𝑩⃗⃗ =(πŸ‘π’ŠΜ‚+πŸ—π’‹Μ‚βˆ’πŸ”π’ŒΜ‚)

(iii) π‘ͺβƒ—βƒ— =(πŸ‘π’ŠΜ‚βˆ’πŸ“π’‹Μ‚+π’ŒΜ‚)

(iv) 𝑫⃗⃗ =(πŸ–π’ŠΜ‚βˆ’πŸ’π’‹Μ‚+πŸ–π’ŒΜ‚)


1
Expert's answer
2021-03-08T08:32:24-0500

The magnitude of vector is:


"a = \sqrt{a_x^2 + a_y^2 + a_z^2}"

where "a_x, a_y, a_z" are coordinates of the vector. Thus, for the given vectors obtain:


"A = \sqrt{2^2 + 3^2 + 6^2} = 7\\
B = \sqrt{3^2 + 9^2 + (-6)^2} \approx 11.2\\
C = \sqrt{3^2 + (-5)^2 + 1^2} \approx 5.9\\
D = \sqrt{8^2 + (-4)^2 + 8^2} = 12"

Answer. 7, 11.2, 5.9, 12.